acetic acid is an important ingredient of vinegar. a sample of 50.0 ml of a commercial vinegar is titrated against a 1.00 M NaOH solution. What is the molarity of acetic acid present in the vinegar if 5.75 ml of the base are needed for the titration?
Answer #1

Consider reaction , CH3COOH (aq) + NaOH (aq)phpDcZmHv.png CH3COONa (aq) + H2O (l)

From reaction, stoichiometric ratio = No. of moles of acid / No. of moles of base = 1 /1 = 1

We have correlation, M acid\times V acid = M base\times V base\times Stoichiometric ratio

i e M CH3COOH\times V CH3COOH = M NaOH\times V NaOH\times Stoichiometric ratio

\therefore M CH3COOH\times 50.0 ml = 1.00 M \times 5.75 ml \times 1

M CH3COOH = 1.00 M \times 5.75 ml \times 1 / 50.0 ml = 0.115 M

ANSWER : Molarity of acetic acid present in the vinegar = 0.115 M